Solve using the multiplication method.Please also do a check.
-1/6x = -9/10
these are fractions and i do not no what i am doing.Help me please.

Answer by stanbon(75887)

-1/6x = -9/10
Cross-multiply to get:
-1*10 = 6x*-9
-10 = -54x
--------------
Divide both sides by -54 to get:
x = 10/54
x = 5/27
===========
Cheers,
Stan H.